OpenStudy (anonymous):

Rational Expression 3P-3/P÷6P-6/3P^2

Nnesha (nnesha):

factor !!!!

Nnesha (nnesha):

\[\huge\rm \frac{3P-3}{P}÷\frac{6P-6}{3P^2}\] is it your question right ?

OpenStudy (anonymous):

@Nnesha yes

OpenStudy (anonymous):

you can flip the right fraction

OpenStudy (anonymous):

and change to multiplication

Nnesha (nnesha):

alright factor (3p-3) what is common in these both terms ? factor (6p-6) and like molly say z that you can change division to multiplication example \[\large\rm \frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

@Nnesha ok than what's netx step

Nnesha (nnesha):

\cancel

Nnesha (nnesha):

after that you can cancel some stuffffffs

OpenStudy (anonymous):

@Nneshais is it 3p/2?

Nnesha (nnesha):

would you like to explain how did you get that answer ? 3p-p what is common factor ?? :-)